If a=c/2 and 3a=7b , find c in terms of b.
Answer by edjones(8007)

a=c/2
c=2a
.
3a=7b
3a * 2/3 = 7b * 2/3
2a=14b/3
c=14b/3 because c=2a
.
Ed